Let A(3,2,-1), B(5,0,-2), C(3,3,0) and D(4,-6,8) be four points in the plane,Donate the plane containing A, B, and C.Use vector methods to solve the following.
(a)Find a unit vector perpendicular to the plane.
(b)WITHOUT finding the equation of the plane , calculate the shortest distance between D and the plane
1
Expert's answer
2020-03-03T16:37:31-0500



Let "\alpha" be the plane containing points A, B, and C. The vector "\vec{n}" is a unit vector perpendicular to the plane. "\overrightarrow{AB} = (2,-2,-1), \overrightarrow{AC} = (0,1,1), \overrightarrow{AD} = (1,-8,9).\\"

a) The unit vector "\vec{n}" can be found as a cross-product of vectors "\overrightarrow{AC}" and "\overrightarrow{AB}" normalized by its length.

"\vec{n} =\dfrac{\overrightarrow{AB} \times \overrightarrow{AC} }{|\overrightarrow{AB} \times \overrightarrow{AC}|}\\" .


"\overrightarrow{AB} \times \overrightarrow{AC} = \begin{vmatrix} \mathbf i & \mathbf j & \mathbf k \\ 2 & -2 & -1 \\ 0 & 1 & 1 \end{vmatrix} = (-1, 2, 2).\\"

"|\overrightarrow{AC} \times \overrightarrow{AB}| = \sqrt{(-1)^2 + (2)^2 + (2)^2} = 3.\\"

Thus, "\vec{n} = \dfrac{1}{3}(-1, 2, 2).\\"


b) The shortest distance "DD'" between D and the plane can be found as a projection of vector "\overrightarrow{AD}" onto the "\vec{n}" direction. Namely,

"DD'=\overrightarrow{AD} \cdot \vec{n}.\\"

"DD' = \dfrac{1}{3}( 1\cdot(-1) + (-8)\cdot2 + 9\cdot 2) = \dfrac{1}{3}"


Answer. a) "\vec{n} = \dfrac{1}{3}(-1, 2, 2)" , b) "DD'= \dfrac{1}{3}" .